# Taming Toxicity or Fueling It? The Great Ban’s Role in Shifting Toxic User Behavior and Engagement

### v1

@@ 4.3 RQ3: Behavioral dynamics of extremely toxic users
Subreddits participation. Initially, we explore the subreddit participation habits of these users as a function of their change in toxicity. This analysis aims to ascertain whether each of the most toxic users engaged in a few or many of the banned subreddits. Figure 5 suggests that only a small number of users participated in multiple subreddits. Instead, the vast majority of users participated only in a few subreddits, typically ranging between one and three, consistently with the echo chamber theory [34]. Moreover, user toxicity changes do not depend on the number of subreddits in which the user participated.
@@
Activity. Next, we look into the relationship between changes in toxicity and activity. We measure user activity as the number of posted comments. Conjointly studying user activity and toxicity is important since toxic users pose a meaningful threat only when high toxicity coincides with elevated activity, which holds the potential to cause a substantial impact on the platform. This joint analysis also enables the identification of those users who are both highly toxic and remarkably active, allowing to prioritize moderation interventions towards them [35]. The left-most plot of Figure 6 presents a density scatterplot of the changes in toxicity and activity for each considered user. The plot area is divided into four quadrants, corresponding to the combinations of increase/decrease in toxicity/activity. As shown, the bulk of the distribution lies close to the plot origin, meaning that the ban caused minor changes in both toxicity and activity. Moreover, those users who experienced marked changes in toxicity only had negligible changes in activity, and vice versa. The insets in Figure 6 show the number of users laying in each quadrant of the scatterplots. The most frequent reaction to the ban was a reduction in both toxicity and activity (38% users lay in Q3). This is also evident in the left-most plot of Figure 7, which pictures the relationship between mean toxicity and user activity, after the ban. It shows how the vast majority of users who exhibit low levels of toxicity are moderately active. The opposite behavior is instead the least frequent (16% users are in Q1). Overall this analysis suggests that the resentful users might have had a limited impact on the platform due to their low activity.
@@ 5 Discussion
Our study provides insights into the consequences of The Great Ban, a crucial example of deplatforming that affected over 2,000 subreddits. We highlight that 15.6% of the users involved abandoned the platform, however those who remained reduced their average toxicity by 4.1%. At the same time, however, approximately 5% of all users significantly escalated their toxicity. The presence of these so-called resentful users was observed across many of the analyzed subreddits. However, only 16% of users increased both their toxicity and activity levels, and just 21% of users who became more toxic also received positive feedback from other users. Our findings provide new insights regarding the impact of The Great Ban and, in general, on the heterogeneous and unintended negative reactions to moderation interventions [8]. Additionally, they shed light on several aspects regarding the design and implementation of effective moderation strategies, underlining the inherent challenges in moderating online platforms.
@@
Effectiveness of the moderation. In the current literature, the effectiveness of content moderation has typically been evaluated by examining the changes in these interventions in terms of activity and toxicity  [15, 6]. In this context, our study showed that a significant portion of toxic users left the platform, while those who remained exhibited a modest decrease in toxicity along with a notable decline in activity. Apparently, these results suggest successful moderation. However, it is crucial to carefully examine these findings by considering potential unintended consequences. For instance, the deplatforming of toxic users may move the issue somewhere else rather than solve it. As highlighted by recent studies, moderated users might migrate to other platforms [7], possibly indicating that their toxic interactions may have been displaced rather than mitigated. Moreover, migrated users may engage in much more toxic and aggressive behavior in the new platforms [7]. Finally, user abandonment and reduced activity after the ban could represent a significant issue for Reddit, as user engagement and interactions are the main sources of revenue for online platforms [5, 36]. As a result, the presumed effectiveness of The Great Ban in reducing toxicity should be interpreted carefully as it can negatively impact not only the platform’s revenue because of abandonment or reduced engagement, but also outside Reddit because of user migration. Future efforts should focus on finding a balance between developing strategies that effectively reduce toxicity while minimizing user abandonment or decline in activity [36]. In fact, effective moderation should foster healthier online communities while safeguarding platform safety and economic sustainability. Otherwise, platforms may be discouraged from implementing rigorous moderation measures.
@@
Divergent reactions to moderation. Our study showed that, despite a slight decrease in toxicity, a significant minority of users exhibited substantial increases in their toxic behavior. This finding has significant implications for evaluating and designing moderation interventions. First, it underscores the complexity of user responses to content moderation, an area that is still vastly underexplored and warrants further investigation [37]. Additionally, it brings to light the necessity for personalization in the context of content moderation. A broad intervention like The Great Ban – which impacted thousands of subreddits and tens of thousands of users – may not adequately fit the different motivations and personalities of those affected. Here, this was shown by the minority of resentful users who significantly escalated their toxic behavior. Understanding the aspects that lead to these divergent responses is crucial for developing effective moderation interventions. Future research and on-field applications should focus on user profiling, by taking into account individual personal traits, past behavior, and contextual factors to effectively customize moderation interventions [35]. In addition, user migration combined to a high percentage of users displaying increased toxicity raises concerns about the potential radicalization effects of moderation and its inadvertent role in amplifying echo chambers [7]. This indicates that some users may not respond positively to specific moderation interventions, potentially resulting in more radical behaviors. Our findings emphasize the careful balance needed in content moderation, which not only should locally reduce toxicity but also should prevent unintended consequences that may lead to polarization or radicalization within specific user groups.
@@
Impact of resentful users. The finding that the vast majority of resentful users – who exhibited increased toxicity following the ban – did not concurrently escalate their activity offers a novel perspective on the little-studied interplay of toxicity and user activity. Contrary to the suspicion that highly toxic users might also be highly active, our results underscore a limited overlap between these two dimensions of user behavior. This suggests that a considerable portion of users expressing heightened toxicity post-ban did not translate their resentment into increased participation on the platform. In turn, this result partly mitigates the previous concerns about the divergent and adverse reactions to The Great Ban. Furthermore, the observation that users who intensified their toxic behaviors did not receive heightened positive social feedback from their peers resonates with some recent research on the drivers of online hateful and toxic speech. The latter posits that online toxicity may be driven by the pursuit of social approval rather than by a desire to harm [20]. To this end, our results demonstrate that the Reddit community did not endorse or support the increased toxicity exhibited by these users. According to the aforementioned theories, toxic users who do not receive the expected levels of social approval should decrease their toxicity over time [20]. However, we measured overall stationary toxicity trends over the course of the 7 months after the ban. Our results thus apparently challenge the social approval theory, in support of competing theories according to which at least a subset of toxic and hateful users seek amusement by harassing others, as in trolling behavior [22, 38]. On the one hand, these results call for additional research to identify the types and characteristics of toxic users. On the other hand, they reinforce the need for nuanced and diversified moderation strategies that are capable of addressing the complexity and multiplicity of online user behavior [39].
@@
Limitations and Future works. In this study, we make use of a large dataset consisting of comments posted by 34K Reddit users during 7 months before and after The Great Ban took place. This means that our findings may be limited to Reddit and the effects of The Great Ban itself. Moreover, online platforms are dynamic and characterized by frequent changes in user behavior, community norms, and platform policies. Our study takes into account a specific timeframe and while it covers a relatively extended period, it limits the generalizability of our findings to different periods of time and somewhat constrains our ability to fully account for long-term or evolving trends. In addition, our evaluation of the toxicity scores could be biased because of the use of Detoxify, a machine learning model. The observational nature of our work constitutes another limitation, as it limits our ability to consider external events that might have independently influenced user behavior outside of The Great Ban. To conclude, data on user demographics, motivations, and contextual factors is missing from our dataset. Having this type of information could have offered a more refined interpretation of our results. Future research could look more into the unexplored connections between user characteristics and moderation outcomes, serving as an initial step towards the implementation of more targeted and personalized moderation interventions [35, 36].
@@ 6 Conclusions
Use of AI
The authors used an LLM to proofread the article prior to submission.

### v3

@@ 4.3 RQ3: Behavioral dynamics of extremely toxic users
Subreddits participation. Initially, we explore the subreddit participation habits of these users as a function of their change in toxicity. This analysis aims to ascertain whether each of the most toxic users engaged in a few or many of the banned subreddits. Figure 5 suggests that only a small number of users participated in multiple subreddits. Instead, the vast majority of users participated only in a few subreddits, typically ranging between one and three, consistently with the echo chamber theory (Cinelli et al., 2021; Valensise et al., 2023). Moreover, user toxicity changes do not depend on the number of subreddits in which the user participated.
@@
Activity. Next, we look into the relationship between changes in toxicity and activity. We measure user activity as the number of posted comments. Conjointly studying user activity and toxicity is important since toxic users pose a meaningful threat only when high toxicity coincides with elevated activity, which holds the potential to cause a substantial impact on the platform. This joint analysis also enables the identification of those users who are both highly toxic and remarkably active, allowing to prioritize moderation interventions towards them (Cresci et al., 2022). The left-most plot of Figure 6 presents a density scatterplot of the changes in toxicity and activity for each considered user. The plot area is divided into four quadrants, corresponding to the combinations of increase/decrease in toxicity/activity. As shown, the bulk of the distribution lies close to the plot origin, meaning that the ban caused minor changes in both toxicity and activity. Moreover, those users who experienced marked changes in toxicity only had negligible changes in activity, and vice versa. The insets in Figure 6 show the number of users laying in each quadrant of the scatterplots. The most frequent reaction to the ban was a reduction in both toxicity and activity (38% users lay in Q3). This is also evident in the left-most plot of Figure 7, which pictures the relationship between mean toxicity and user activity, after the ban. It shows how the vast majority of users who exhibit low levels of toxicity are moderately active. The opposite behavior is instead the least frequent (16% users are in Q1). Overall this analysis suggests that the resentful users might have had a limited impact on the platform due to their low activity.
@@ 5 Discussion
Our study provides insights into the consequences of The Great Ban, a crucial example of deplatforming that affected over 2,000 subreddits. We highlight that 15.6% of the users involved abandoned the platform, however those who remained reduced their average toxicity by 4.1%. At the same time, however, approximately 5% of all users significantly escalated their toxicity. The presence of these so-called resentful users was observed across many of the analyzed subreddits. However, only 16% of users increased both their toxicity and activity levels, and just 21% of users who became more toxic also received positive feedback from other users. Our findings provide new insights regarding the impact of The Great Ban and, in general, on the heterogeneous and unintended negative reactions to moderation interventions (Trujillo and Cresci, 2023). Additionally, they shed light on several aspects regarding the design and implementation of effective moderation strategies, underlining the inherent challenges in moderating online platforms.
@@
Effectiveness of the moderation. In the current literature, the effectiveness of content moderation has typically been evaluated by examining the changes in these interventions in terms of activity and toxicity  (Chandrasekharan et al., 2017; Jhaver et al., 2021). In this context, our study showed that a significant portion of toxic users left the platform, while those who remained exhibited a modest decrease in toxicity along with a notable decline in activity. Apparently, these results suggest successful moderation. However, it is crucial to carefully examine these findings by considering potential unintended consequences. For instance, the deplatforming of toxic users may move the issue somewhere else rather than solve it. As highlighted by recent studies, moderated users might migrate to other platforms (Horta Ribeiro et al., 2021), possibly indicating that their toxic interactions may have been displaced rather than mitigated. Moreover, migrated users may engage in much more toxic and aggressive behavior in the new platforms (Horta Ribeiro et al., 2021). Finally, user abandonment and reduced activity after the ban could represent a significant issue for Reddit, as user engagement and interactions are the main sources of revenue for online platforms (Trujillo and Cresci, 2022; Tessa et al., 2024). As a result, the presumed effectiveness of The Great Ban in reducing toxicity should be interpreted carefully as it can negatively impact not only the platform’s revenue because of abandonment or reduced engagement, but also outside Reddit because of user migration. Future efforts should focus on finding a balance between developing strategies that effectively reduce toxicity while minimizing user abandonment or decline in activity (Tessa et al., 2024). In fact, effective moderation should foster healthier online communities while safeguarding platform safety and economic sustainability. Otherwise, platforms may be discouraged from implementing rigorous moderation measures.
@@
Divergent reactions to moderation. Our study showed that, despite a slight decrease in toxicity, a significant minority of users exhibited substantial increases in their toxic behavior. This finding has significant implications for evaluating and designing moderation interventions. First, it underscores the complexity of user responses to content moderation, an area that is still vastly underexplored and warrants further investigation (Jhaver et al., 2023). Additionally, it brings to light the necessity for personalization in the context of content moderation. A broad intervention like The Great Ban – which impacted thousands of subreddits and tens of thousands of users – may not adequately fit the different motivations and personalities of those affected. Here, this was shown by the minority of resentful users who significantly escalated their toxic behavior. Understanding the aspects that lead to these divergent responses is crucial for developing effective moderation interventions. Future research and on-field applications should focus on user profiling, by taking into account individual personal traits, past behavior, and contextual factors to effectively customize moderation interventions (Cresci et al., 2022; Annamoradnejad et al., 2022). In addition, user migration combined to a high percentage of users displaying increased toxicity raises concerns about the potential radicalization effects of moderation and its inadvertent role in amplifying echo chambers (Horta Ribeiro et al., 2021). This indicates that some users may not respond positively to specific moderation interventions, potentially resulting in more radical behaviors. Our findings emphasize the careful balance needed in content moderation, which not only should locally reduce toxicity but also should prevent unintended consequences that may lead to polarization or radicalization within specific user groups.
@@
Impact of resentful users. The finding that the vast majority of resentful users – who exhibited increased toxicity following the ban – did not concurrently escalate their activity offers a novel perspective on the little-studied interplay of toxicity and user activity. Contrary to the suspicion that highly toxic users might also be highly active, our results underscore a limited overlap between these two dimensions of user behavior. This suggests that a considerable portion of users expressing heightened toxicity post-ban did not translate their resentment into increased participation on the platform. In turn, this result partly mitigates the previous concerns about the divergent and adverse reactions to The Great Ban. Furthermore, the observation that users who intensified their toxic behaviors did not receive heightened positive social feedback from their peers resonates with some recent research on the drivers of online hateful and toxic speech. The latter posits that online toxicity may be driven by the pursuit of social approval rather than by a desire to harm (Jiang et al., 2023). To this end, our results demonstrate that the Reddit community did not endorse or support the increased toxicity exhibited by these users. According to the aforementioned theories, toxic users who do not receive the expected levels of social approval should decrease their toxicity over time (Jiang et al., 2023). However, we measured overall stationary toxicity trends over the course of the 7 months after the ban. Our results thus apparently challenge the social approval theory, in support of competing theories according to which at least a subset of toxic and hateful users seek amusement by harassing others, as in trolling behavior (Chandrasekharan et al., 2022; Uyheng et al., 2022). On the one hand, these results call for additional research to identify the types and characteristics of toxic users. On the other hand, they reinforce the need for nuanced and diversified moderation strategies that are capable of addressing the complexity and multiplicity of online user behavior.
@@
Limitations and Future works. In this study, we make use of a large dataset consisting of comments posted by 34K Reddit users during 7 months before and after The Great Ban took place. This means that our findings may be limited to Reddit and the effects of The Great Ban itself. Moreover, online platforms are dynamic and characterized by frequent changes in user behavior, community norms, and platform policies. Our study takes into account a specific timeframe and while it covers a relatively extended period, it limits the generalizability of our findings to different periods of time and somewhat constrains our ability to fully account for long-term or evolving trends. In addition, our evaluation of the toxicity scores could be biased because of the use of Detoxify, a machine learning model. The observational nature of our work constitutes another limitation, as it limits our ability to consider external events that might have independently influenced user behavior outside of The Great Ban. To conclude, data on user demographics, motivations, and contextual factors is missing from our dataset. Having this type of information could have offered a more refined interpretation of our results. Future research could look more into the unexplored connections between user characteristics and moderation outcomes, serving as an initial step towards the implementation of more targeted and personalized moderation interventions (Cresci et al., 2022; Tessa et al., 2024).
@@ Acknowledgments
Use of AI
During the preparation of this work the author(s) used ChatGPT in order to proofread the article prior to its submission. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.